The lifetime of bulbs produced by Tariqs lightening company follows a normal distribution with a mean of 3462 hours and a standard deviation 1070 hours. What is the probability that randomly selected bulb will have a lifetime of at most 2800?

1
Expert's answer
2022-01-30T14:29:10-0500

What is the probability that a randomly chosen light bulb will have a lifetime of at most 2800 hours?

"μ = 3462 \; hours \\

σ = 1070\;hours \\

P(X<2800) \\

= P(Z< \frac{2800-3462}{1070} )\\

= P(Z<-0.6187) \\

=0.26806 \\"